OpenStudy (lena772):

Mike poured 15 cups of water into a rectangular container measuring 10 inches by 6 inches by 9 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points) Part C: What was the height of the water in the container? Show your work. (4 points) Measurement is given as length by width by height.

OpenStudy (bibby):

I've been told my handwriting is ugly so I'll try explaining unit cancellation one more time \[\frac{ 1 inch }{ 2.54 cm }\] if we want to convert 30 cm to inches we can just multiply and cancel the units out in our head\[30 cm * \frac{ 1 inch }{ 2.54 cm } = \frac{ 30 }{ 2.54 } = 11.811 inches\]

OpenStudy (bibby):

In part B we know 15 cups were poured in. Converting 15 cups into cubic inches using the formula they gave us, [1 cup = 14.44 cubic inches] \[15cups*\frac{ 14.44 inches^3 }{ 1 c up} = 15*14.44 cubic inches = 216.6 inches \]

OpenStudy (the_fizicx99):

@bibby where did you get 30cm

OpenStudy (bibby):

That was just a number I pulled out of nowhere for unit cancellation. The way to approach THIS problem is a) get them into one unit and multiply b) copy whatever I just did

OpenStudy (the_fizicx99):

Part A: V = L * W * H V = 10" * 6" * 9 cm (convert 9cm to inches = 9 * 2.54 = 22.86) V = 10 * 6 * 22.86 V = 1371.6 Now all measurement are in the same unit. Part B: 1 cup is = 14.44 cubic inches So, 14.44 * 15 = 216.6 216.6 cubic inches of water have been poured into the container. Part C: The maximum volume of water that the rectangular container can hold is V = 1371.6 Mike poured 216.6 cubic inches into the pan. 216.6 = 10 * 6 * H 216.6 / 60 = 3.61 Idk I think I forgot something ?

OpenStudy (the_fizicx99):

Think its my volume and height

OpenStudy (bibby):

high five brother you just embarrassed this old man excuse me while I take a seat and cry

OpenStudy (the_fizicx99):

Wacha mean :o

OpenStudy (anonymous):

(convert 9cm to inches = 9 * 2.54 = 22.86) This is incorrect

OpenStudy (bibby):

9/2.54 owned idiot nerd

OpenStudy (anonymous):

Like it doesnt even make sense logically

OpenStudy (the_fizicx99):

Oh!! Knew something was wrong u.u

OpenStudy (the_fizicx99):

@bibby why would you divide 9/2.54??

OpenStudy (lena772):

To get the inches

OpenStudy (lena772):

2.54 cm in an inch

OpenStudy (bibby):

Yeah that's why I was spending time talking about cancelling the units in your mind. You have a ratio\[ \frac{ 1 inch }{ 2.54cm }\]

OpenStudy (anonymous):

Its all abt proportions 1 inch : 2.54 cm x inch: 9 cm \(\large \frac{1}{x}=\frac{2.54}{9}\) \( 9=2.54x\) \(\large \frac{9}{2.54}=x\)

OpenStudy (anonymous):

Ok my method is confusing -_-

OpenStudy (lena772):

Idk who is right here....

OpenStudy (bibby):

I'm using a chemistry method of unit cancellation lol

OpenStudy (anonymous):

Im using what I learned in 8th grade lol

OpenStudy (lena772):

so fizic is wrong?

OpenStudy (bibby):

ok I'll just start from the beginning and outline what you're supposed to do at each step. a: convert 9 cm into inches (divide by 2.54) multiply that number, 10, and 6 to get the volume of the whole container. B: convert 15 cups into cubic inches by multiplying by 14.44 That is the amount poured into the container Part C We have all the dimensions but the height [number from part b] = 10*6*x

OpenStudy (lena772):

Thank you~